OpenStudy (anonymous):

How would I find the last two digits of 79^79 ? I know that 79^79 mod 100 = 79^39 mod 100 since: 79 mod phi(100) = 79 mod 40 = 39. But after that I'm lost. Help would be appreciated.

OpenStudy (shubhamsrg):

Brilliant.org eh ? ;)

OpenStudy (anonymous):

Yes ;) It was a bad example sorry. I've read the blog but I still don't get it. Take something else. Last two digits of 13^13 ?

OpenStudy (shubhamsrg):

I'll hint you 79^5 = -1 mod100

OpenStudy (shubhamsrg):

Well I am not an expert at modular arithmetic, but my approach for this one was to find a power of 79 close to 100, manually, 79^5 was one

OpenStudy (anonymous):

Sounds reasonable, I'll try it. Thanks

terenzreignz (terenzreignz):

Well, well, something looks eerily familiar :D How'd you know 79^5 was -1(mod 100), @shubhamsrg XD

OpenStudy (shubhamsrg):

As I said, I manually calculated it, used a calculator

terenzreignz (terenzreignz):

sounds masochistic :D

OpenStudy (shubhamsrg):

hmm, in case you have a shorter way ?

terenzreignz (terenzreignz):

Reminds me of this one http://openstudy.com/study#/updates/510a825ae4b070d859bf144e

terenzreignz (terenzreignz):

No, I was just reminiscing... good times XD

OpenStudy (anonymous):

Okey the n^k = -1 mod 100 seems to be important but why and how?

OpenStudy (shubhamsrg):

hmm, maybe I see something 79 = 80-1 79^79 = (80-1)^79 All terms on expansion, of RHS will be divisible by 100, except , C(79,0) * (-1) + C(79,1) * (80) = (79*80)-1 omg it was so easy! :O

terenzreignz (terenzreignz):

Not important, maybe lucky :D Because if a = b mod 100 a^n = b^n mod 100 And I don't know about you, but raising -1 (or 1) for that matter to exponents seems like a relatively simple task :D

OpenStudy (anonymous):

aha now i see. 79^5^8 = -1^8 mod 100 79^40 = 1 mod 100 cool

OpenStudy (shubhamsrg):

I wasted a lot of time on this ques when I was myself attempting it a few hours before. Now it seems a lot easier 79= 80-1 Thats it.